# Ambient Diffusion Posterior Sampling: Solving Inverse Problems with Diffusion Models trained on Corrupted Data

### v1

@@ 2.3 Ambient Diffusion Posterior Sampling
$\mathrm{d}{\bm{x}}=-2\dot{\sigma}_{t}\sigma_{t}\left(\underbrace{\frac{\mathbb{E}[{\bm{x}}_{0}|{\bm{y}}_{t,\mathrm{train}},A_{\mathrm{train}}]-{\bm{x}}_{t}}{\sigma_{t}}}_{\mathrm{Ambient\ Score}}+\gamma_{t}\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}_{\mathrm{inf}}|{\bm{x}}_{0}=\mathbb{E}[{\bm{x}}_{0}|{\bm{y}}_{t,\mathrm{train}},A_{\mathrm{train}}])\right)\mathrm{d}t+g(t)\mathrm{d}{\bm{w}},$ (2.12)
for a fixed $A_{\mathrm{train}}\sim p(A_{\mathrm{train}})$. Comparing this to the DPS update rule (E.q. 2.3), all the $\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]$ terms have been replaced with their ambient counterparts, i.e. with $\mathbb{E}[{\bm{x}}_{0}|{\bm{y}}_{t,\mathrm{train}},A_{\mathrm{train}}]$. We remark that similar to DPS, the proposed algorithm is an approximation to sampling from the true posterior distribution $\mathbb{E}[{\bm{x}}_{0}|{\bm{y}}_{\inf}]$. We term our approximate sampling algorithm for solving inverse problems with diffusion models learned from corrupted data Ambient DPS (A-DPS).
@@ 3.1 Ambient Diffusion with Pre-trained Models on Datasets of Natural Images
In this section, we evaluate the performance of Ambient DPS, which uses diffusion models trained on corrupted data, and we compare it to DPS, which uses diffusion models trained on clean data. For our experiments, we use the models from the Ambient Diffusion [17] that are trained on randomly inpainted data with different erasure probabilities. Specifically, for AFHQ we use the Ambient Models with erasure probability $p\in\{0.2,0.4,0.6,0.8\}$ and for Celeb-A we use the pre-trained models with $p\in\{0.6,0.8,0.9\}$.
We underline that all the Ambient Models have worse performance for unconditional generation compared to the models trained with clean data (i.e. the models trained with $p=0.0$). The goal of this work is to explore the conditional generation performance of Ambient Models, where the conditioning is in the measurements ${\bm{y}}_{\mathrm{inf}}$, and compare it with models trained on uncorrupted data. To ensure that Ambient Models do not have an unfair advantage, we test only on restoration tasks that are different from the ones encountered in their training. Specifically, we use models trained on random inpainting and we evaluate on Gaussian Compressed Sensing [4] and Super Resolution.
Hyperparameters.
The only tunable parameters for DPS (Eq. 2.3) and Ambient DPS (Eq. 2.12) are in the scheduling of the magnitude of the measurements likelihood term. In all the experiments in the DPS paper, this term is kept constant throughout the diffusion sampling trajectory and the authors recommend selecting a value in the range between $[0.1,10]$. We follow this recommendation and we keep this term constant. The value of the step size for each model is selected with a hyperparameter search in the recommended range. For all our experiments, we follow exactly the DPS implementation provided in the official code repository of the paper. The other parameter that impacts performance is the number of steps we are going to run each algorithm for, i.e. the discretization level of the SDEs of Equations 2.3, 2.12. Typically, the higher the number of steps the better the performance since the discretization error decreases [10, 11]. For the performance results, we run each method for several steps $\in\{50,100,150,200,250,300\}$, and we report the best result among them.
@@
Figure 4: Speed performance plots for AFHQ.
@@ Results.
Figure 2 presents Gaussian Compressed Sensing reconstruction results (i.e. reconstructing a signal from Gaussian random projections). We show MSE and LPIPS performance metrics for the AFHQ dataset as we vary the number of measurements. The results are given for models that are trained with inpainted images at different levels of corruption, indicated by the erasure probability $p$. As shown in the Figure, the model trained with clean data outperforms the models trained with corrupted data when the number of measurements is high. However, as we reduce the number of measurements, Ambient Models outperform the models trained with clean data in the very low measurements regime. To the best of our knowledge, there is no known theoretical argument that explains this performance cross-over, and understanding this further is an interesting research direction. Similar results are presented in Figure 3 for the task of super-resolution at AFHQ. The model trained on clean data ($p=0$) slightly outperforms the Ambient Models in both LPIPS and MSE for reconstructing a 2×2$\times$2 × downsampled image, as expected. Yet, as the resolution decreases, there is again a cross-over in performance and models trained on corrupted data start to outperform the models trained on uncorrupted data. We include results for LPIPS and MSE for Compressed Sensing and Downsampling in FFHQ and Celeb-A in the Appendix (Figs 10, 11, 12, 13).
@@
Finally, we examine how the number of sampling steps affects the performance. The MSE results for Compressed Sensing with $4000$ measurements on AFHQ are shown in Figure 4. As shown, the higher the erasure probability $p$ during training, the better the Compressed Sensing performance of the model for low Number of Function Evaluations (NFEs). Models trained with higher corruption are faster since they require fewer steps for the same performance. For increased NFEs, the models that are trained on clean(er) data finally outperform. This result is consistent across different datasets (AFHQ, FFHQ, CelebA), reconstruction tasks (Compressed Sensing, Downsampling), and metrics (MSE, LPIPS) (Figures 14, 15, 16, 17, 18 in the Appendix).
3.2 Ambient MRI Diffusion Models
In the previous section, we used pre-trained diffusion models on datasets of natural images. In this section, we extend Ambient Diffusion to the multi-coil Fourier subsampled MRI setting (as detailed in Section 2.2) and we train our own MRI models from scratch.
Table 1: Unconditional Sampling FID Scores for models trained on MRI data at different acceleration factors.
Inference/Training Training Data R Reconstruction Method FID
EDM $1$ MVUE $10.4082$
EDM $2$ L1-Wavelet $18.5534$
EDM $4$ L1-Wavelet $27.6405$
EDM $6$ L1-Wavelet $51.4272$
EDM $8$ L1-Wavelet $102.984$
NCSNV2 $1$ MVUE $13.2006$
Ambient $2$ Adjoint $30.3425$
Ambient $4$ Adjoint $32.3101$
Ambient $6$ Adjoint $31.5033$
Ambient $8$ Adjoint $48.1525$
Dataset.
For the dataset preparation, we follow standard practices in the MRI literature. Specifically, we download the FastMRI dataset, randomly select 2,000 T2-Weighted scans, and pick 5 center slices from each scan. This gives us a dataset of 10,000 T2-weighted brain scans ($k$-space measurements). We pass each multi-channel $k$-space sample through a noise pre-whitening filter to transform the noise into standard white Gaussian. Then, we normalize each $k$-space sample by the absolute value of $99$-th percentile of the root sum-of-squares reconstruction of the autocalibration region (24x24 pixels from the center). Given the pre-processed fully sampled $k$-space samples, we estimate sensitivity maps using the ESPIRiT calibration method [44]. These measurements serve as a fully sampled reference dataset (i.e., $R=1$).
Sampling masks.
We note that, based on the Ambient Diffusion theory, the extra corruption is a free parameter that can be chosen to be minimal, and perhaps, we could have chosen to corrupt the available data less aggressively. Due to the lack of computational resources, we did not ablate the degree of further corruption and we leave this as future work.
Comparison Models.
We train one model for each acceleration factor $R\in\{1,2,4,6,8\}$. The $R=1$ model is trained directly on clean data (no extra corruption) based on the EDM approach [26]. The $R>1$ models are trained with our modified Ambient Diffusion framework. While we focus on EDM for fully sampled training, we also train an NCSNV2 model for $R=1$ following the approach in [25]. As an additional baseline, we also train models with EDM after L1-Wavelet compressed sensing reconstruction of the training set at each acceleration factor [33]. We used the BART implementation [45, 6] and searched for the best regularization parameter over the training set. Specifically, we used a regularization weighting of 0.001 with 100 iterations. We point the interested reader to Section E in the Appendix (and to our open-sourced code) for all the training hyperparameters.
Figure 5: Prior samples from Ambient Diffusion trained with under sampled data at $R=2,4,6,8$ (columns $1-4$), and EDM trained with fully sampled data (column $5$).
Figure 6: Posterior reconstructions using A-DPS (columns $2-5$) and FS-DPS (column $6$). Rows $1$ and $3$ show reconstructions at $R=4$ and $R=8$, respectively, while rows $2$ and $4$ display their difference images on a 10×10$\times$10 × scale.
Unconditional generation evaluation.
We first evaluate the unconditional generation performance of each model. We show prior samples for the Ambient Diffusion models and EDM at $R=1$ in Figure 5. Additional samples for all trained models are included in the Appendix, Figure 19. Similar to the results in [17, ], we observe that the Ambient Diffusion samples are qualitatively similar to the EDM models at low acceleration, and become slightly blurry at higher accelerations. Notably, the samples generated by Ambient Diffusion at $R=8$ have no residual aliasing artifacts, in contrast to the samples generated from L1-Wavelet reconstructions (as seen in the Appendix).
To further quantify unconditional sample quality, we follow the approach in [5] and calculate FID scores from 100 samples using a pre-trained VGG network. Table 1, shows the FID scores for the different diffusion models. While there is an increase in FID for the Ambient models, we see that those trained at higher acceleration factors outperform the models trained on L1-Wavelet reconstructions.
3.3 MRI Reconstruction
After verifying that our trained models have learned the underlying distribution reasonably well, we want to evaluate how they perform in the task of accelerated image reconstruction. We start by listing the details of our models and the considered baselines.
Ambient Diffusion Posterior Sampling (A-DPS)
We implemented our method by following Eq. 2.12 for $500$ steps with a likelihood weighting $\gamma_{t}=\frac{1}{\left\|{\bm{y}}-A\mathbb{E}[{\bm{x}}_{0}|{\bm{y}}_{t,\mathrm{train}},A_{\mathrm{train}}]\right\|_{2}}$.
Ambient One Step (A-OS)
As mentioned previously, our method also admits a one-step solution. This is outlined in section 2.4 by noting that we train our model to estimate $\mathbb{E}[{\bm{x}}_{0}|\tilde{\bm{y}}_{\mathrm{train}},\tilde{A}_{\mathrm{train}}]$. Thus we can use our models in a one-step fashion. The performance of the one-step prediction should only be expected to be good in-distribution, i.e. when the model is evaluated at the same acceleration factor as the one used during training.
L1-Wavelet Compressed Sensing
We use the same L1-Wavelet reconstruction described previously as a standalone non-deep learning comparison.
Self-Supervised learning via Data Under-sampling (SSDU) [47]
In a supervised setting, end-to-end image restoration networks can be trained by passing measurements through a network like MoDL [2] or VarNet [22] and taking a loss directly on the output of the network with the ground truth image. In the self-supervised setting, however, this is not possible. SSDU, trains the end-to-end restoration network by splitting the available measurement set ($\Omega$) of each sample $y^{\Omega}_{i}$ into training sets ($\Theta$) and loss sets ($\Lambda$) or measurements $y^{\Theta}_{i}$ and $y^{\Lambda}_{i}$ respectively where $\Omega=\Theta\cup\Lambda$. Where the reconstruction network is given access to measurements in the training set to obtain an estimated image $\hat{{\bm{x}}}^{\Theta}_{i}={\bm{h}}_{\theta}(y^{\Theta}_{i})$ and the loss is then defined over the loss measurements as
$L({\bm{y}}^{\Lambda}_{i},\hat{{\bm{x}}}^{\Theta}_{i})=\frac{\left\|{\bm{y}}^{\Lambda}_{i}-A_{\Lambda}\hat{{\bm{x}}}^{\Theta}_{i}\right\|_{1}}{\left\|{\bm{y}}^{\Lambda}_{i}\right\|_{1}}+\frac{\left\|{\bm{y}}^{\Lambda}_{i}-A_{\Lambda}\hat{{\bm{x}}}^{\Theta}_{i}\right\|_{2}}{\left\|{\bm{y}}^{\Lambda}_{i}\right\|_{2}}$ (3.1)
where $A_{\Lambda}$ is the forward operator with ones in the selection mask $P$ at only the measurement locations in the loss set and zeros elsewhere. In this work, we selected ${\bm{h}}_{\theta}(\cdot)$ to have the MoDL architecture [2]. We trained individual models at four different acceleration levels ($R=2,4,6,8$). Each model was trained for ten epochs using the same training set ($10,000$ slices) as previously described. We used uniform random sampling to separate the training and loss measurement groups for each sample and found that a split of $\rho=\frac{|\Lambda|}{|\Omega|}=0.2$ provided the best performance and thus was used for reporting all subsequent metrics.
Figure 7: NRMSE performance metrics at $R=2,4,6,8$ for all methods.
Figure 8: PSNR performance metrics at $R=2,4,6,8$ for all methods.
Figure 9: SSIM performance metrics at $R=2,4,6,8$ for all methods.
MoDL Supervised (MoDL-Sup)
To provide an upper bound on performance for supervised end-to-end methods we also trained the MoDL architecture in a supervised fashion [2]. Again, we trained individual models at four different acceleration levels ($R=2,4,6,8$). Each model was trained using the same $10,000$ slices as above. However, for these models, we used the normalized root mean squared error (NRMSE) as the loss function:
$L({\bm{x}}_{i},\hat{{\bm{x}}}_{i})=\frac{\left\|{\bm{x}}_{i}-\hat{{\bm{x}}}_{i}\right\|_{2}}{\left\|{\bm{x}}_{i}\right\|_{2}}$ (3.2)
Diffusion Posterior Sampling (DPS)
For models trained on clean data, we can use the Diffusion Posterior Sampling (DPS) [12] algorithm. We implemented DPS with a variety of trained diffusion models. Specifically, we implemented the DPS algorithm with ($\mathbf{1}$) a Diffusion model trained with fully sampled ground truth images (FS-DPS), ($\mathbf{2-5}$) four different diffusion models trained on the L1-Wavelet reconstructions of data from measurement sets at $R=2,4,6,8$ (L1-DPS). Inference was run using the update rule of Eq. 2.3, for a total of $500$ steps and with $\gamma_{t}=\frac{1}{\left\|{\bm{y}}-A\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]\right\|_{2}}$.
Annealed Langevin Dynamics (ALD)
Another popular method for solving inverse problems via score-based priors is with ALD [25]. Here we trained a score-based model using the same fully sampled $10,000$ samples as above and ran inference for $1,300$ steps. This method acts as another fully sampled comparison to our technique and thus serves a similar purpose as FS-DPS.
Results
Quantitative metrics for the listed methods are included in Figs. 7, 8, 9 under the NRMSE, PSNR and SSIM metrics respectively. We observe that at low acceleration levels ($R=2,4$) FS-DPS outperforms most methods including all A-DPS models. This is not unexpected, since FS-DPS uses models trained on clean data. However, as the validation acceleration increases to higher ratios $R$, A-DPS begins to outperform all models, including the ones trained on fully sampled data such as FS-DPS and ALD. This trend is consistent with the results that we observed in Section 3.1 for Ambient Models trained on natural images.
To get a better sense of the performance of the different methods, we visualize reconstructions for a subset of the methods listed at various accelerations in Fig. 6. Here, we see that at lower accelerations ($R=4$), just as the metrics suggest, FS-DPS outperforms A-DPS. However, at higher acceleration ($R=8$) we can see that FS-DPS introduces significant artifacts into the reconstruction while our A-DPS reconstructions maintain a high degree of visual fidelity.
4 Conclusions
We presented a simple framework based on DPS for solving inverse problems with Ambient Diffusion models. We showed that diffusion models trained on missing data are state-of-the-art inverse problem solvers for high corruption levels even if the forward process at inference time is different from the one used during training. Our framework fully unlocks the potential of Ambient Diffusion models that are critical in applications where access to full data is impossible or undesirable.
5 Acknowledgements
This research has been supported by NSF Grants CCF 1763702, NSF CCF-2239687 (CAREER), AF 1901292, CNS 2148141, Tripods CCF 1934932, NSF AI Institute for Foundations of Machine Learning (IFML) 2019844, Oracle for Research Fellowship, Google Research Scholars Award, the Texas Advanced Computing Center (TACC) and research gifts by Western Digital, WNCG IAP, UT Austin Machine Learning Lab (MLL), Cisco and the Archie Straiton Endowed Faculty Fellowship. Giannis Daras has been supported by the Onassis Fellowship (Scholarship ID: F ZS 012-1/2022-2023), the Bodossaki Fellowship and the Leventis Fellowship.
Appendix A Theorem Proof

### v2

@@ 2.3 Ambient Diffusion Posterior Sampling
$\mathrm{d}{\bm{x}}=-2\dot{\sigma}_{t}\sigma_{t}\left(\underbrace{\frac{\mathbb{E}[{\bm{x}}_{0}|\tilde{\bm{y}}_{t,\mathrm{train}},\tilde{A}_{\mathrm{train}}]-{\bm{x}}_{t}}{\sigma_{t}}}_{\mathrm{Ambient\ Score}}+\gamma_{t}\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}_{\mathrm{inf}}|{\bm{x}}_{0}=\mathbb{E}[{\bm{x}}_{0}|\tilde{\bm{y}}_{t,\mathrm{train}},\tilde{A}_{\mathrm{train}}])\right)\mathrm{d}t+g(t)\mathrm{d}{\bm{w}},$ (2.11)
for a fixed $\tilde{A}_{\mathrm{train}}\sim p(\tilde{A}_{\mathrm{train}})$  [^1] The Ambient MRI models take as input $\tilde{P}$ instead of $\tilde{A}$, as in Equation 2.10.. Comparing this to the DPS update rule (E.q. 2.3), all the $\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]$ terms have been replaced with their ambient counterparts, i.e. with $\mathbb{E}[{\bm{x}}_{0}|\tilde{\bm{y}}_{t,\mathrm{train}},\tilde{A}_{\mathrm{train}}]$. The latter is approximated by a neural network that is trained with Ambient Diffusion.
We term our approximate sampling algorithm for solving inverse problems with diffusion models learned from corrupted data Ambient DPS (A-DPS). We remark that similar to DPS, the proposed algorithm is an approximation to sampling from the true posterior distribution. Similarly to Theorem 1 of the DPS paper, for a given measurement ${\bm{y}}_{\mathrm{inf}}$ at noise level $\sigma_{{\bm{y}}}$ and a given matrix $A_{\mathrm{inf}}$, one can upper-bound the Jensen gap of A-DPS as follows:
$\mathcal{J}\leq\frac{n}{\sqrt{2\pi\sigma_{{\bm{y}}}^{2}}}\exp(-1/2\sigma_{{\bm{y}}}^{2})||A_{\mathrm{inf}}||\int_{{\bm{x}}_{0}}||{\bm{x}}_{0}-\hat{\bm{x}}_{0}||p({\bm{x}}_{0}|{\bm{x}}_{t})\mathrm{d}{\bm{x}}_{0},$ (2.12)
where $\hat{\bm{x}}_{0}=\mathbb{E}[{\bm{x}}_{0}|\tilde{A}_{\mathrm{train}}{\bm{x}}_{t},\tilde{A}_{\mathrm{train}}]$, for a fixed $\tilde{A}_{\mathrm{train}}$ sampled from $p(\tilde{A}_{\mathrm{train}})$. We omit the proof of this proposition since it follows the exact same steps as the proof of Theorem 1 in the DPS paper.
@@
Table 1: Unconditional and Conditional Sampling FID for models trained on MRI data at different acceleration factors.
Training Data Training Method Unconditional FID $\downarrow$
$R=1$ EDM 10.41
NCSNV2 $13.20$
$R=2$ L1-EDM 18.55
Ambient Diffusion $30.34$
$R=4$ L1-EDM 27.64
Ambient Diffusion $32.31$
$R=6$ L1-EDM $51.43$
Ambient Diffusion 31.50
$R=8$ L1-EDM $102.98$
Ambient Diffusion 48.15
Training Data Inference Data Reconstruction Method Conditional FID $\downarrow$
$R=1$ $R=8$ FS-DPS $4.16$
$R=2$ $R=8$ A-DPS (our method) 0.71
$R=4$ $0.87$
$R=6$ $1.29$
$R=8$ $2.00$
@@ 3 Experiments
3.1 Ambient MRI Diffusion Models
In this section, we extend Ambient Diffusion to the multi-coil Fourier subsampled MRI setting (as detailed in Section 2.2) and we train our own MRI models from scratch. We give details about our dataset preparation in Section B in the Appendix.
Sampling masks.
Comparison Models.
We train one model for each acceleration factor $R\in\{1,2,4,6,8\}$. The $R=1$ model is trained on clean data (no extra corruption) based on the EDM approach (Karras et al., 2022). The $R>1$ models are trained with our modified Ambient Diffusion framework. While we focus on EDM for fully sampled training, we also train an NCSNV2 model for $R=1$ following the approach in Jalal et al. (2021). As baselines, we also train EDM models after L1-Wavelet compressed sensing reconstruction (L1-EDM) of the training set at each acceleration factor (Lustig et al., 2007).
Unconditional generation evaluation.
We evaluate the unconditional generation performance of each model and show prior samples for Ambient Diffusion and EDM models at $R=1$ in Figure 2. Similar to the results in Daras et al. (2023b), we observe that the Ambient Diffusion samples are qualitatively similar to EDM models at low acceleration, and become slightly blurry at higher accelerations. Notably, the samples generated by Ambient Diffusion at $R=8$ have no residual aliasing artifacts, in contrast to the samples generated from L1-Wavelet reconstructions.
To quantify unconditional sample quality, we follow the approach in Bendel et al. (2022) and calculate FID scores from 100 samples using a pre-trained VGG network. Table 1, shows FID scores for the diffusion models. While there is an increase in FID for the Ambient models, we see that those trained at higher acceleration factors outperform L1-EDM models trained on L1-Wavelet reconstructions.
Table 2: Accelerated MRI reconstruction performance evaluation using Normalized Root Mean Squared Error (NRMSE $\downarrow$), Structural Similarity Index Measure (SSIM $\uparrow$), and Peak Signal-to-Noise Ratio (PSNR $\uparrow$) averaged across 100 test samples.
3.2 Accelerated MRI Reconstruction
We now evaluate the accelerated reconstruction performance of Ambient Diffusion models trained on subsampled MRI data. Our method, A-DPS was implemented by following Eq. 2.11 for $500$ steps with likelihood weighting $\gamma_{t}=\frac{1}{\left\|{\bm{y}}-A\mathbb{E}[{\bm{x}}_{0}|{\bm{y}}_{t,\mathrm{train}},A_{\mathrm{train}}]\right\|_{2}}$. We consider the following baselines:
  • Fully-Sampled Diffusion Posterior Sampling (FS-DPS). For models trained on clean data, we can use the Diffusion Posterior Sampling (DPS) (Chung et al., 2023) algorithm. Specifically, we implemented the DPS algorithm with a Diffusion model trained with fully sampled ground truth images (FS-DPS). Inference was run using the update rule of Eq. 2.3, for a total of $500$ steps and with $\gamma_{t}=\frac{1}{\left\|{\bm{y}}-A\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]\right\|_{2}}$.
  • Fully-Sampled Annealed Langevin Dynamics (FS-ALD). A popular method for solving inverse problems via score-based priors is ALD (Jalal et al., 2021). We train a score-based model using the same fully sampled $10,000$ samples as above and run inference for $1,300$ steps. This method acts as another fully sampled comparison to our technique.
  • L1-Wavelet Compressed Sensing (L1-CS). We use the same L1-Wavelet reconstruction described previously as a standalone non-deep learning comparison.
  • Model Based Deep Learning Supervised (MoDL-Sup). We use a supervised end-to-end model using the MoDL architecture in a supervised fashion (Aggarwal et al., 2019) for each acceleration factor. For inference, we pass the undersampled data through the trained MoDL network. Further details of this baseline are mentioned in Section C in the appendix.
  • Ambient Annealed Langevin Dynamics (A-ALD). We utilize our Ambient Diffusion MRI models with the inverse problem solver ALD (Jalal et al., 2021) for $1,300$ steps.
  • Ambient One Step (A-OS). Our method also admits a one-step solution. This is outlined in section 2.4 by noting that we train our model to estimate $\mathbb{E}[{\bm{x}}_{0}|\tilde{\bm{y}}_{\mathrm{train}},\tilde{A}_{\mathrm{train}}]$. Thus we can use our models in a one-step fashion. The performance of the one-step prediction should only be expected to be good in-distribution, i.e. when the model is evaluated at the same acceleration factor as the one during training.
@@ 3.2 Accelerated MRI Reconstruction
  • L1-Diffusion Posterior Sampling (L1-DPS). We utilize the diffusion models trained on the L1-Wavelet reconstructions of data from measurement sets at $R=2,4,6,8$. For reconstruction, we use DPS for a total of $500$ steps with $\gamma_{t}=\frac{1}{\left\|{\bm{y}}-A\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]\right\|_{2}}$.
  • Self-Supervised learning via Data Under-sampling (SSDU). We train SSDU (Yaman et al., 2020) models using the end-to-end restoration network by splitting the available measurements ($\Omega$) of each sample $y^{\Omega}_{i}$ into training sets ($\Theta$) and loss sets ($\Lambda$) or measurements $y^{\Theta}_{i}$ and $y^{\Lambda}_{i}$. Further SSDU training details are given in Section C.
  • 1D-Partitioned SSDU. We train 1D-Partitioned SSDU (Millard & Chiew, 2023) which stems from the SSDU framework and introduces partitioning of the sampling set to ensure each subset has a distribution similar to the original mask. The loss function, network architecture, and training hyperparameters were set according to Millard & Chiew (2023).
Results
Quantitative metrics for the listed methods are included in Table 2 under the NRMSE, SSIM and PSNR metrics respectively. Furthermore, we report feature-based metrics including LPIPS and DISTS in Table 3. At low acceleration levels ($R=2,4$) FS-DPS outperforms most methods, including A-DPS. However, as the validation acceleration increases to higher ratios $R$, A-DPS outperforms other models, including those trained on fully sampled data.
We also trained an Ambient Diffusion model on undersampled abdominal MRI scans, where no ground truth data is available (Zhang et al., 2013). We show visual examples from the prior and posterior sampling experiments in the Appendix (Figs 7, 8).
We visualize reconstructions for a subset of the methods listed at various accelerations in Fig. 3. Here, we see that FS-DPS outperforms A-DPS at lower accelerations ($R=4$). However, at higher acceleration ($R=8$) we can see that FS-DPS introduces significant artifacts into the reconstruction while our A-DPS reconstructions maintain a high degree of visual fidelity. We also report the distribution metrics (Conditional FID) between samples from the posterior and the ground-truth distribution for A-DPS and FS-DPS methods in Table 1. Similar to our findings from posterior reconstruction metrics, we find that in the higher acceleration ($R=8$) regime, A-DPS outperforms FS-DPS. We also include graphical trends for NRMSE, SSIM, and PSNR for A-DPS, A-OS, FS-DPS, L1-DPS, FS-ALD, L1-CS, MoDL-Sup, and SSDU in the Appendix (Figs 9, 10, 11, 12, 13).
We use the models from the Ambient Diffusion (Daras et al., 2023b) that are trained on randomly inpainted data with different erasure probabilities. Specifically, for AFHQ we use the Ambient Models with erasure probability $p\in\{0.2,0.4,0.6,0.8\}$ and for Celeb-A we use the pre-trained models with $p\in\{0.6,0.8,0.9\}$. We underline that all the Ambient Models have worse performance for unconditional generation than those trained with clean data (i.e. the models trained with $p=0.0$). This work aims to explore the conditional generation performance of Ambient Models, where the conditioning is in the measurements ${\bm{y}}_{\mathrm{inf}}$, and compare it with models trained on uncorrupted data. To ensure that Ambient Models do not have an unfair advantage, we test only on restoration tasks that differ from those encountered in their training. Specifically, we use models trained on random inpainting and we evaluate Gaussian Compressed Sensing (Baraniuk, 2007) and Super Resolution.
@@ Results.
In Figure 4 we show visual examples from our experiment on the noisy super-resolution task for natural images using a model: (1) trained with fully-sampled data (FS-DPS), (2) trained with randomly inpainted data at $p=0.6$ (A-DPS). Figure 5 presents Gaussian Compressed Sensing reconstruction results (i.e. reconstructing a signal from Gaussian random projections). We show MSE and LPIPS performance metrics for the AFHQ dataset across varying number of measurements. The results are given for models that are trained with inpainted images at different levels of corruption, indicated by the erasure probability $p$. The model trained with clean data outperforms the models trained with corrupted data when the number of measurements is high. However, as we reduce the number of measurements, Ambient Models outperform the models trained with clean data in the very low measurements regime. To the best of our knowledge, there is no known theoretical argument that explains this performance cross-over, and understanding this further is an interesting research direction. Similar results are presented in Figure 6 for the task of super-resolution at AFHQ. The model trained on clean data ($p=0$) slightly outperforms the Ambient Models in both LPIPS and MSE for reconstructing a 2×2$\times$2 × downsampled image, as expected. Yet, as the resolution decreases, there is again a cross-over in performance and models trained on corrupted data start to outperform the models trained on uncorrupted data. We include results for LPIPS and MSE for Compressed Sensing and Downsampling in FFHQ and Celeb-A in the Appendix (Figs 15, 16, 17, 18).
@@
We compute MSE in other inverse problem settings for FFHQ: Box Inpainting, Additive Gaussian Noise, and Super Resolution (with Gaussian Noise $\sigma=0.05$). We find that trends in this experiment are consistent with previous findings. The results across multiple downsampling factors and noise levels in the Appendix (Figs 19, 20, 21). Finally, we examine how the number of sampling steps affects the performance. The MSE results for Compressed Sensing with $4000$ measurements on AFHQ are shown in Figure 14. As shown, the higher the erasure probability $p$ during training, the better the Compressed Sensing performance of the model for low Number of Function Evaluations (NFEs). Models trained with higher corruption are faster since they require fewer steps for similar performance. For increased NFEs, the models trained on clean(er) data outperform. This result is consistent across different datasets (AFHQ, FFHQ, CelebA), reconstruction tasks (Compressed Sensing, Downsampling), and metrics (MSE, LPIPS) (Figs 22, 23, 24, 25, 26 in the Appendix).
4 Conclusion
We present Ambient Diffusion Posterior Sampling (A-DPS), a simple framework based on DPS for solving inverse problems with Ambient Diffusion models. We show that Ambient Diffusion models trained on corrupted data can be better suited for handling ill-posed inverse problems under severe corruption. By being exposed to corrupted training data, A-DPS exhibits robust priors that generalize better to high-acceleration regimes, unlike FS-DPS, which may overfit clean data and fail to adapt effectively to severe undersampling. Our framework fully unlocks the potential of Ambient Diffusion models that are critical in applications where access to full data is impossible or undesirable.
Appendix A Theoretical Results
A.1 Theory Proof